# Cascaded Multi-Scale Attention for Enhanced Multi-Scale Feature Extraction and Interaction with Low-Resolution Images

### v1

@@ 3.4.2 Design of a Block
We choose $3\times 3$ depth-wise convolution for $\mathrm{PE}$. We incorporate a train-time overparametrization strategy (Chu et al., 2021; Vasu et al., 2023a), aiming at a better trade-off between inference accuracy and computational cost. Specifically, our network features three parallel branches in this section during the training phase: a $3\times 3$ depth-wise convolution, a point-wise convolution, and batch normalization. These branches are reparametrized into a single depth-wise convolution at the testing phase. Following the depth-wise convolution, we utilize the GeLU activation function.
@@
where BN denotes Batch Normalization. Its output is fed into a feed-forward network (FFN), following the standard design of ViTs (Dosovitskiy et al., 2021) as
@@
The patch embedding layer comprises reparameterizable depth-wise and point-wise convolution layers, that use train-time overparameterization and these additional branches are removed by structural reparameterization at inference (Ding et al., 2021a, b).
3.4.3 Other Components
The Stem block and the patch embedding layers situated between the stages are each implemented as a convolutional layer.Specifically, we adopt a single pair consisting of a depth-wise convolution followed by a point-wise convolution for each, aiming to boost computational efficiency. For implementation, we adopt a train-time overparametrization strategy (Vasu et al., 2023b), similar to the one used for positional embedding, to achieve a favorable balance between accuracy and efficiency.
Table 3: Results on landmark-free head pose estimation with the BIWI Fanelli et al. (2013) and AFLW2000 Zhu et al. (2016) datasets. Models are trained on the 300W-LP Zhu et al. (2016) dataset. means the values reported in the respective papers.
Model Input Params BIWI AFLW2000
@@ 3.4.3 Other Components
Hopenet($\alpha$=2)  (Ruiz et al., 2018) 224$\times$224 23.0 5.17 6.98 3.39 5.18 6.47 6.56 5.44 6.16
Shao et al. (Shao et al., 2019) 224$\times$224 24.6 4.59 7.25 6.15 6.00 5.07 6.37 4.99 5.48
WHENet  (Zhou and Gregson, 2020) 224$\times$224 4.4 3.99 4.39 3.06 3.81 5.11 6.24 4.92 5.42
TokenHPE-v2  (Zhang et al., 2023a) 224$\times$224 85.9 3.95 4.51 2.71 3.72 4.36 5.54 4.08 4.66
QuatNet  (Hsu et al., 2019) 227$\times$227 6.8 4.01 5.49 2.94 4.15 3.97 5.62 3.92 4.50
FDN  (Zhang et al., 2020b) 224$\times$224 5.8 4.52 4.70 2.56 3.93 3.78 5.61 3.88 4.42
Li et al. (Li et al., 2022a) 224$\times$224 0.9 3.59 3.94 2.68 3.40 3.36 5.05 3.56 3.99
EVA-GCN  (Xin et al., 2021) 256$\times$256 3.3 4.46 5.34 4.11 4.64 4.01 4.78 2.98 3.92
Li et al. (Li et al., 2022a) 56$\times$56 0.9 4.85 5.92 3.29 4.69 4.48 6.55 4.80 5.28
FSA-Net  (Yang et al., 2019) 64$\times$64 5.1 4.27 4.96 2.76 4.00 4.50 6.08 4.64 5.07
TriNet  (Cao et al., 2021) 64$\times$64 26.0 4.11 4.76 3.05 3.97 4.04 5.77 4.20 4.67
CMSA-B 64$\times$64 5.4 4.47 5.29 3.15 4.30 4.08 5.30 4.06 4.48
CMSA-L 64$\times$64 7.1 4.49 4.63 2.85 3.99 4.15 5.10 3.89 4.38
TriNet  (Cao et al., 2021) 32$\times$32 26.0 4.96 7.02 3.75 5.24 6.79 9.10 7.54 7.81
Li et al. (Li et al., 2022a) 28$\times$28 0.9 6.49 9.12 4.71 6.77 6.52 8.56 7.19 7.42
CMSA-S 32$\times$32 4.0 4.81 4.85 3.24 4.33 4.01 5.65 4.12 4.59
CMSA-B 32$\times$32 5.4 5.28 5.25 3.13 4.55 3.88 5.58 4.07 4.51
CMSA-L 32$\times$32 7.1 4.97 5.36 3.12 4.48 3.84 5.54 3.99 4.46
4 Experimental Results
@@
4.1 Human Pose Estimation
4.1.1 Dataset
We utilize the COCO 2017 dataset  (Lin et al., 2014), comprising training, validation, and test sets with 118,000, 5,000, and 20,000 images, respectively. We adopt the bottom-up setting employed in many previous studies, in which the individual persons’ bounding boxes are provided [^4] We use the official Github repository of HRNet (Sun et al., 2019) as a framework for the experiments: https://github.com/leoxiaobin/deep-high-resolution-net.pytorch. . Each person is labeled with 17 keypoints, and the models are tasked with predicting these keypoints’ coordinates based on the cropped and resized image from its bounding box in the input image. The dataset contains 250,000 individuals, with their heights in the original images ranging from $32$ pixels to over $128$ pixels throughout the dataset. These images are resized to a standard size before being fed into a model and we adjust this size to manipulate the input resolution.
4.1.2 Training
We consider the three variants with different model size, CMSA-S, -B, and -L, as shown in Table 1. We trained our CMSAs from scratch for 210 epochs using a mean squared error loss function, following the recipe from DeiT (Touvron et al., 2021), which is summarized as follows. We employed the AdamW (Loshchilov and Hutter, 2019) optimizer and applied a linear warm-up for the first five epochs. The initial learning rate was set to $2.5\times 10^{-3}$, and the minimum value was set to $10^{-5}$. The weight decay and momentum were set to $5\times 10^{-2}$ and $0.9$, respectively, and a batch size of 32 was used. Data augmentation includes random rotation, random scale, flipping, and random erasing during training. For comparison, we consider several state-of-the-art methods. For them, we employ their official repositories [^5] UDP: https://github.com/HuangJunJie2017/UDP-Pose,
@@ 4.1.2 Training
PCT: https://github.com/Gengzigang/PCT.
P2T: https://github.com/yuhuan-wu/P2T,
SG-Former:https://github.com/OliverRensu/SG-Former,
Shunted: https://github.com/OliverRensu/Shunted-Transformer,
SMT: https://github.com/AFeng-x/SMT,
SBCFormer: https://github.com/xyongLu/SBCFormer.
to train and evaluate them using smaller-sized images, which are not tested in their respective papers.
Fig. 3: Average Precision vs. Parameters. All models are trained for human pose estimation with input size $32\times 32$.
4.1.3 Results
@@
Table 2 presents the results at various input resolutions. Following prior studies, we employ the standard evaluation metrics based on object keypoint similarity (OKS) to gauge the accuracy of the models. Our CMSAs demonstrate superior performance across all tested resolutions, notably surpassing competitors by a wider margin at lower input resolutions [^6] It is noteworthy that they tend to show slightly lower AR scores, it is attributable to the choice of learning rate schedulers; when trained with the same step scheduler (StepLRScheduler in PyTorch) as others including ViT-B, the models (CMSA-B* and -L* in Table 2) yield superior AR scores.. This underscores our model’s efficacy in delivering high performance for low-resolution images.
@@
Figure 3 presents the performance of methods using typical backbones with multi-scale representation for human pose estimation on $32\times 32$ images. Our CMSAs consistently outperform the comparison methods, demonstrating superior multi-scale representation abilities on low-resolution images.
4.2 Head Pose Estimation
4.2.1 Datasets
We apply our model to head pose estimation in a setting that does not rely on facial landmarks 777 FSA-Net: https://github.com/shamangary/FSA-Net,
TriNet: https://github.com/anArkitek/TriNet_WACV2021.
. Models are tasked with predicting the pitch, roll, and yaw angles of a person’s head from an appropriately cropped image. Before inputted to the models, these images are resized to a standard resolution. We adjust this resolution to manage the size of the input images effectively. Building on prior research, we train our model and the SOTA models on the 300W-LP dataset and assess their performance using both the BIWI and AFLW2000 datasets. BIWI dataset has roughly 15,000 frames generated from RGB-D videos captured by a Kinect device for different subjects and head poses. 300W-LP dataset contains 61,225 images and further expanded 122,450 samples with flipping. Most of the faces in the BIWI dataset are small angles (yaw: ±$\pm 75^{\circ}$, pitch: ±$\pm 60^{\circ}$, roll: ±$\pm 50^{\circ}$), while 300W-LP extends the original 300-W dataset to a large pose ([$-90^{\circ}$, $90^{\circ}$]).The AFLW2000 dataset is from the first 2,000 images of the AFLW dataset (Koestinger et al., 2011). Following the previous studies (Ruiz et al., 2018; Yang et al., 2019), we discard 31 images with angles outside of [$-99^{\circ}$, $99^{\circ}$] in the AFLW2000 dataset.
4.2.2 Training
During training, we applied binned classification and soft stage-wise regression to estimate head pose from a single RGB image, as in (Yang et al., 2019). Our method was trained for 100 epochs on the 300W-LP dataset following the recipe from DeiT. The training utilized AdamW optimizer with consine learning rate scheduling and an initial learning rate $2.5\times 10^{-3}$. Batch size is set to 96. During training and evaluation, all images are cropped around the face to include the whole head, resized to $32\times 32$, and normalized using the mean and standard deviation from ImageNet dataset (Deng et al., 2009).
4.2.3 Results
Table 3 shows the results of ours and the state-of-the-art methods, where the standard evaluation metric, the mean absolute error (MAE), is adopted. At low-resolution settings of $64\times 64$ and $32\times 32$, our CMSAs can obtain state-of-the-art performance for the AFLW2000 dataset. Especially, CMSA-L achieves an MAE of 4.46 on $32\times 32$ images from AFLW2000, surpassing the performance of most typical models for higher image resolutions. For low-resolution images from BIWI dataset, our CMSA-L consistently achieves state-of-the-art performance.
4.3 Image Classification with CIFAR-10/100
4.3.1 Training
We train our models from scratch for 300 epochs, following DeiT’s (Touvron et al., 2021) recipe, with the same data augmentation, cross-entropy loss function, and AdamW optimizer. The initial learning rate was set to $2.5\times 10^{-3}$, and the batch size was 128. For data augmentation, we apply conventional techniques such as random cropping, random horizontal flipping, mixup, random erasing, and label smoothing. We adhere to the original authors’ code when training other models. All models undergo training under uniform conditions with respect to the number of epochs and data augmentation strategies during both training and testing phases.
@@ 4.3.1 Training
Fig. 4: Accuracy vs. Parameters. All models are trained for CIFAR-10 classification with input size $32\times 32$.

### v3

@@ III-D2 Design of a Block
We choose $3\times 3$ depth-wise convolution for $\mathrm{PE}$. We incorporate a train-time overparametrization strategy [40, 33], aiming at a better trade-off between inference accuracy and computational cost. Specifically, our network features three parallel branches in this section during the training phase: a $3\times 3$ depth-wise convolution, a point-wise convolution, and batch normalization. These branches are reparameterized into a single depth-wise convolution at the testing phase. We apply GeLU activation after the depth-wise convolution.
@@
where BN denotes Batch Normalization. Its output is fed into a feed-forward network (FFN), following the standard design of ViTs [26] as
@@
The patch embedding layer comprises reparameterizable depth-wise and point-wise convolution layers, which use train-time overparameterization, and these additional branches are removed by structural reparameterization at inference [41].
III-D3 Other Components
The Stem block and patch embedding layers between stages are implemented as a convolutional layer. Specifically, we adopt a single pair consisting of a depth-wise convolution followed by a point-wise convolution for each, aiming to boost computational efficiency. For implementation, we adopt a train-time overparametrization strategy [42], similar to the one used for positional embedding, to achieve a favorable balance between accuracy and efficiency.
TABLE II: Results on bottom-up human pose estimation with the COCO 2017 val dataset. All FLOPs are calculated at the respective single-scale. The speed of all methods is recorded on a single RTX 3090 GPU with a batch size of 32. means the results reported in the respective papers. * indicates models trained with a different learning rate scheduler.
Model Backbone Input Params FLOPs Speed COCO 2017 val
@@ III-D3 Other Components
SimpleBaseline [43] ResNet-50 $128\times 96$ 34.0 2.3 697.1 59.3 85.5 67.4 57.8 63.8 66.6
HRNet-W32 [6] HRNet-W32 $128\times 96$ 28.5 1.9 318.1 66.9 88.7 76.3 64.6 72.3 73.7
HRNet-W48 [6] HRNet-W48 $128\times 96$ 63.6 3.9 316.3 68.0 88.9 77.4 65.7 73.7 74.7
DARK [44] HRNet-W32 $128\times 96$ 28.5 1.9 257.4 70.7 88.9 78.4 67.9 76.6 76.7
UDP [45] HRNet-W32 $128\times 96$ 28.6 1.9 307.8 69.5 91.3 77.7 67.9 72.2 73.7
HRFormer-B [2] HRFormer-B $128\times 96$ 43.2 3.7 176.6 66.2 88.2 76.1 64.1 71.8 73.0
ViTPose-B [46] ViT-B $128\times 96$ 86.0 17.0 190.6 67.8 88.3 76.2 65.9 73.3 74.6
PCT-B [47] Swin-B $128\times 96$ 87.0 11.6 137.2 66.4 90.5 75.7 65.1 68.7 70.1
CMSA-B CMSA-B $128\times 96$ 5.6 4.1 761.0 70.3 91.3 77.8 67.9 77.4 74.6
CMSA-L CMSA-L $128\times 96$ 7.3 5.1 728.2 71.5 91.4 79.2 69.0 75.4 75.5
HRNet-W32 [6] HRNet-W32 $64\times 48$ 28.5 1.9 321.9 58.3 86.2 67.1 57.4 62.3 66.2
CAL [21] HRNet-W48 $64\times 48$ 110.3 5.3 - 61.5 88.1 68.7 60.7 63.5 66.3
UDP [45] HRNet-W32 $64\times 48$ 28.6 1.9 311.4 62.8 86.4 70.9 61.3 67.8 70.4
HRFormer-B [2] HRFormer-B $64\times 48$ 43.2 3.7 176.0 60.9 87.1 70.6 60.0 65.2 68.7
ViTPose-B [46] ViT-B $64\times 48$ 86.0 16.9 191.9 63.8 86.5 72.6 62.5 68.7 71.4
PCT-B [47] Swin-B $64\times 48$ 87.0 3.9 174.7 62.0 90.5 71.9 61.4 63.8 66.7
CMSA-B CMSA-B $64\times 48$ 5.6 3.7 809.7 64.7 87.6 72.3 62.9 69.7 71.7
CMSA-L CMSA-L $64\times 48$ 7.3 4.6 785.8 65.9 88.1 74.5 64.0 70.2 72.7
CMSA-B CMSA-B $64\times 48$ 5.6 3.7 809.7 65.2 91.4 75.3 63.0 68.7 69.3
CMSA-L CMSA-L $64\times 48$ 7.3 4.6 785.8 66.0 91.5 76.4 63.6 69.5 70.0
HRNet-W32 [6] HRNet-W32 $32\times 24$ 28.5 1.9 326.8 42.6 79.6 41.5 43.4 44.6 52.7
UDP [45] HRNet-W32 $32\times 24$ 28.6 1.9 312.5 51.9 81.7 56.8 51.5 55.9 61.4
HRFormer-B [2] HRFormer-B $32\times 24$ 43.2 3.7 177.5 45.4 81.8 46.2 45.9 48.0 55.3
ViTPose-B [46] ViT-B $32\times 24$ 86.0 16.9 193.5 51.2 81.0 56.1 50.7 55.0 60.4
PCT-B [47] Swin-B $32\times 24$ 87.0 3.9 178.6 44.4 84.6 41.8 45.3 43.7 50.7
CMSA-S CMSA-S $32\times 24$ 4.1 0.7 1078.8 52.3 84.0 57.3 51.6 53.9 57.2
CMSA-B CMSA-B $32\times 24$ 5.6 0.9 1005.4 53.5 84.9 58.1 53.2 54.7 58.5
CMSA-L CMSA-L $32\times 24$ 7.3 1.1 986.9 56.4 87.0 62.6 55.6 58.2 61.1
IV Experimental Results
@@
IV-A Human Pose Estimation
IV-A1 Dataset
We use the COCO 2017 dataset [48], which includes 118,000 training and 5,000 validation, and 20,000 test images. Following the common bottom-up setting [^4] We use the official Github repository of HRNet [6] as a framework for the experiments: https://github.com/leoxiaobin/deep-high-resolution-net.pytorch. , individual persons’ bounding boxes are provided. Each person is labeled with 17 keypoints. The dataset contains 250,000 individuals with heights ranging from 32 to over 128 pixels. These images are resized to a standard size before being fed into a model, and we adjust this size to manipulate the input resolution.
IV-A2 Training
We consider three variants with different model sizes, CMSA-S, -B, and -L, as shown in Table I. Models are trained from scratch for 210 epochs with mean squared error loss, following the DeiT recipe [49]. Training uses AdamW [50] with linear warm-up for five epochs, an initial learning rate of $2.5\times 10^{-3}$ (minimum $10^{-5}$), weight decay $5\times 10^{-2}$, momentum $0.9$, and batch size 32. Data augmentation includes random rotation, random scale, flipping, and random erasing. For comparison, we evaluate state-of-the-art methods using their official repositories [^5] UDP: https://github.com/HuangJunJie2017/UDP-Pose,
@@ IV-A2 Training
PCT: https://github.com/Gengzigang/PCT,
. They are trained and evaluated on smaller images, which were not tested in the original papers.
IV-A3 Results
@@
Table II presents the results at various input resolutions. Following prior studies, we employ the standard evaluation metrics based on object keypoint similarity (OKS) to gauge the accuracy of the models. Our CMSAs demonstrate superior performance across all tested resolutions, notably surpassing competitors by a wider margin at lower input resolutions [^6] It is noteworthy that they tend to show slightly lower AR scores, it is attributable to the choice of learning rate schedulers; when trained with the same step scheduler (StepLRScheduler in PyTorch) as others including ViT-B, the models (CMSA-B* and -L* in Table II) yield superior AR scores.. This underscores our model’s efficacy in delivering high performance for low-resolution images.
@@
Figure 3 presents the performance of methods using typical backbones with multi-scale representation for human pose estimation on $32\times 24$ images. Our CMSAs consistently outperform the comparison methods, demonstrating superior multi-scale representation abilities on low-resolution images.
@@
Figure 3: Average Precision vs. Parameters. All models are trained for human pose estimation with input size $32\times 24$.
IV-B Head Pose Estimation
IV-B1 Datasets
We apply our model to head pose estimation without relying on facial landmarks, where models predict pitch, roll, and yaw angles directly from cropped images. Models are trained on 300W-LP [51] and tested on BIWI [52] and AFLW2000 [51]. 300W-LP dataset contains 61,225 images (122,450 samples with flipping) covering large poses ($^{\circ}, 9090^{\circ}$]), while BIWI has roughly 15,000 frames with small angles (yaw: $\pm 75^{\circ}$, pitch: $\pm 60^{\circ}$, roll: $\pm 50^{\circ}$). AFLW2000 provides 2,000 images with 3D annotations and 68 landmarks. Following prior work [53, 54], we discard 31 images with angles outside of $^{\circ}, 9999^{\circ}$] in AFLW2000.
IV-B2 Training
We apply binned classification and soft stage-wise regression to estimate head pose from a single RGB image. Models are trained for 100 epochs on the 300W-LP dataset following the recipe from DeiT [49], with AdamW, cosine learning rate scheduling, and an initial learning rate $2.5\times 10^{-3}$. Input images are cropped to include the full head and normalized with ImageNet statistics [55], using a batch size of 96. We use the official repositories for baselines777 FSA-Net: https://github.com/shamangary/FSA-Net,
TriNet: https://github.com/anArkitek/TriNet_WACV2021,
TokenHPE-V2: https://github.com/zc2023/TokenHPE,
6DoF-HPE: https://github.com/Redhwan-A/6DoFHPE,
6DRepNet: https://github.com/thohemp/6DRepNet360.
TABLE III: Results on landmark-free head pose estimation with the BIWI and AFLW2000 datasets. Models are trained on the 300W-LP dataset. All FLOPs are calculated at single-scale. means the results reported in the respective papers.
Model Input Params FLOPs BIWI AFLW2000
Resolution (M) (G) $\downarrow Yaw Pitch Roll MAE$ $\downarrow$
Hopenet($\alpha$=2) [53] $\times224$ 23.9 4.1 5.17 6.98 3.39 5.18 6.47 6.56 5.44 6.16
Shao et al. [56] $\times224$ 24.6 4.1 4.59 7.25 6.15 6.00 5.07 6.37 4.99 5.48
WHENet [57] $\times224$ 4.4 0.4 3.99 4.39 3.06 3.81 5.11 6.24 4.92 5.42
TokenHPE-v2 [58] $\times224$ 86.4 17.0 3.95 4.51 2.71 3.72 4.36 5.54 4.08 4.66
FDN [59] $\times224$ - - 4.52 4.70 2.56 3.93 3.78 5.61 3.88 4.42
Li et al. [60] $\times224$ 0.2 - 3.59 3.94 2.68 3.40 3.36 5.05 3.56 3.99
EVA-GCN [61] $\times256$ $\sim$3.3 - 4.46 5.34 4.11 4.64 4.01 4.78 2.98 3.92
MFDNet [62] $\times224$ - - 3.40 4.68 2.77 3.62 4.30 5.16 3.69 4.38
6DoF-HPE [63] $\times224$ 37.9 8.2 3.91 4.43 2.69 3.68 3.56 4.74 3.35 3.88
6DRepNet [64] $\times224$ 43.8 11.2 3.23 5.32 2.78 3.78 3.27 4.58 2.98 3.61
Li et al. [60] $\times56$ 0.2 - 4.85 5.92 3.29 4.69 4.48 6.55 4.80 5.28
TokenHPE-v2 [58] $\times64$ 85.9 17.0 5.79 5.72 2.98 4.83 4.54 6.43 4.55 5.17
FSA-Net [54] $\times64$ 1.2 - 4.27 4.96 2.76 4.00 4.50 6.08 4.64 5.07
TriNet [65] $\times64$ 26.0 1.1 4.11 4.76 3.05 3.97 4.04 5.77 4.20 4.67
2DHeadPose [66] $\times64$ 24.7 1.6 4.42 5.41 3.37 4.40 4.17 6.61 4.59 5.12
6DoF-HPE [63] $\times64$ 37.9 2.5 4.56 4.79 2.80 4.05 4.80 6.12 4.32 5.08
6DRepNet [63] $\times64$ 43.8 2.9 4.56 4.64 2.85 4.02 4.89 5.87 4.38 5.04
CMSA-B $\times64$ 5.4 4.7 4.47 5.29 3.15 4.30 4.08 5.30 4.06 4.48
CMSA-L $\times64$ 7.1 5.8 4.49 4.63 2.85 3.99 4.15 5.10 3.89 4.38
TriNet [65] $\times32$ 26.0 1.1 4.96 7.02 3.75 5.24 6.79 9.10 7.54 7.81
Li et al. [60] $\times28$ 0.2 - 6.49 9.12 4.71 6.77 6.52 8.56 7.19 7.42
TokenHPE-v2 [58] $\times32$ 85.7 16.9 6.83 6.41 3.76 5.66 5.35 7.34 5.46 6.05
2DHeadPose [66] $\times32$ 24.7 1.5 6.06 7.49 3.73 5.76 5.14 7.55 5.61 6.10
6DoF-HPE [63] $\times32$ 37.9 2.4 5.78 7.80 3.56 5.71 6.23 6.72 5.08 6.01
6DRepNet [63] $\times32$ 43.8 2.8 6.69 6.14 3.69 5.51 6.37 6.85 4.98 6.07
CMSA-S $\times32$ 4.0 0.9 4.81 4.85 3.24 4.33 4.01 5.65 4.12 4.59
CMSA-B $\times32$ 5.4 1.2 5.28 5.25 3.13 4.55 3.88 5.58 4.07 4.51
CMSA-L $\times32$ 7.1 1.5 4.97 5.36 3.12 4.48 3.84 5.54 3.99 4.46
IV-B3 Results
Table III shows the results of ours and the state-of-the-art methods, where the standard evaluation metric, the mean absolute error (MAE), is adopted. At low-resolution settings of $64\times 64$ and $32\times 32$, our CMSAs can obtain state-of-the-art performance for the AFLW2000 dataset. Especially, CMSA-L achieves an MAE of 4.46 on $32\times 32$ images from AFLW2000, surpassing the performance of most typical models for higher image resolutions. For low-resolution images from BIWI dataset, our CMSA-L consistently achieves state-of-the-art performance.
Figure 4: Accuracy vs. Parameters. All models are trained for CIFAR-10 classification with input size $32\times 32$.
IV-C Image Classification with CIFAR-10/100
IV-C1 Training
We train our CMSA and existing models from scratch for 300 epochs on CIFAR-10 and CIFAR-100 at a resolution of $32\times 32$, following DeiT’s [49] recipe, with identical data augmentation, cross-entropy loss function, and AdamW optimizer. The initial learning rate was $2.5\times 10^{-3}$ with a batch size of 128.